OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-8, -10) and is perpendicular to -8x – 3y = -16

OpenStudy (anonymous):

slope intercept form is y=mx+b you need to put the given eq in this form in order to begin

OpenStudy (anonymous):

I don't know what you mean. :/

OpenStudy (anonymous):

put -8x-3y=-16 into slope intercept form

OpenStudy (anonymous):

I really am not sure how to do that.

OpenStudy (anonymous):

can you turn the manipulate the equation algebraically to get y on the left side and x and whatever else on the right side

OpenStudy (cwtan):

The gradient of the equation given: -8x-3y=-16 8x+3y=16 3y=16-8x y=16/3-(8/3)x gradient =-8/3 For two line which is perpendicular,both gradient is multiply each other =-1 m1m2=-1 Therefore gradient for the answer is 3/8 y=(3/8)x+c to find c(y-intercept) sub x= -8 y=-10 into the equation -10=-3+c c=-7 So the equation is y=(3/8)x-7

OpenStudy (anonymous):

thank you!!